# TabMDA: Tabular Manifold Data Augmentation for Any Classifier using Transformers with In-context Subsetting

### v1

@@ Abstract
Tabular data is prevalent in many critical domains, yet it is often challenging to acquire in large quantities. This scarcity usually results in poor performance of machine learning models on such data. Data augmentation, a common strategy for performance improvement in vision and language tasks, typically underperforms for tabular data due to the lack of explicit symmetries in the input space. To overcome this challenge, we introduce TabMDA, a novel method for manifold data augmentation on tabular data. This method utilises a pre-trained in-context model, such as TabPFN, to map the data into a manifold space. TabMDA performs label-invariant transformations by encoding the data multiple times with varied contexts. This process explores the manifold of the underlying in-context models, thereby enlarging the training dataset. TabMDA is a training-free method, making it applicable to any classifier. We evaluate TabMDA on five standard classifiers and observe significant performance improvements across various tabular datasets. Our results demonstrate that TabMDA provides an effective way to leverage information from pre-trained in-context models to enhance the performance of downstream classifiers.
@@ 1 Introduction
To address the challenge of small tabular datasets, we introduce TabMDA, a new training-free method for performing manifold data augmentation (MDA) on tabular data. TabMDA is a general method that works with any downstream classifier. It has two components: first, it leverages an existing pre-trained in-context model, such as TabPFN [17], to embed the data in a manifold. Second, we introduce in-context subsetting (ICS), a novel technique that uses an in-context classifier to facilitate label-invariant transformations directly in the manifold space. ICS encodes the input data multiple times using an in-context encoder, with each iteration using different data subsets as context for the transformer model. This process explores the manifold space learned by the pre-trained encoder (see Figure 1 for an illustration). The downstream classifier is then trained on this expanded manifold dataset, thereby indirectly incorporating the pre-training knowledge into any subsequent classification model. Crucially, our method requires no additional training and is applicable to any classifier. We investigate TabMDA using five common tabular classifiers across 28 different dataset configurations of varying sizes. Our results show that TabMDA substantially improves the performance of various standard tabular classifiers, including tree-based ensembles such as XGBoost, while reducing the variability among their performance. As a result, TabMDA enables explainable classifiers, such as KNN, to become very competitive and sometimes even achieve the highest performance.
@@ 2 Related Work
Transformer-based models for tabular data are typically developed by training from scratch on the target task. Occasionally, an additional pre-training phase on the target task itself is included, unlike LLMs, which use large external corpora. These models introduce various attention mechanisms to handle the tabular data structure. For instance, TUTA [48] introduces a novel attention mechanism that allows the model to attend to the entire table at once, while TAPAS [16] introduces a column-specific attention mechanism that allows the model to attend to the relevant columns in the table. TransTab [49] combines column descriptions and cells as input to a gated Transformer model for feature encoding. SAINT [43] performs attention across both rows and columns to capture the relationships between different cells in the table. More notable examples in this type of models include TabNet [1], TabTransformer [18], and FT-Transformer [14]. Even though transformers can be competitive on large tabular tasks but are often outperformed by well-tuned MLPs [21], especially on small datasets [28]. This suggests that models trained on a single tabular task lack adequate in-context learning (ICL) capabilities for tabular data.
@@
In-context learning (ICL) for tabular data and TabPFN. ICL allows Transformer-based models to adapt to new tasks by conditioning on demonstrations of input-output pairs. Recently, ICL has shown promising results on tabular data. Notably, TabPFN [17], a meta-learned Transformer with ICL capabilities, has demonstrated competitive accuracy compared to tree-based ensembles on small supervised classification tasks. TabPFN employs a Prior-Data Fitted Network (PFN) citepPFN approach, where it is pre-trained on synthetic datasets generated from Structural Causal Models (SCMs) and Bayesian Neural Networks (BNNs), which resembles the causal structure commonly assumed in tabular data. This pretraining enables TabPFN to approximate Bayesian inference in a single forward pass, reducing computational needs and eliminating hyper-parameter tuning. During inference, TabPFN uses the training dataset (context) and new sample features as input, directly outputting the posterior predictive distribution (PPD) for the new samples. This leverages in-context learning to embed new samples without parameter updates, making TabPFN effective on small datasets. Although TabPFN can embed samples into a rich manifold space using its in-context learning abilities, the nature of this embedding space has not been studied. However, we believe these manifolds could facilitate data augmentation for various classifiers.
@@
Tabular Data Augmentation. Tabular data augmentation is less explored due to several challenges, such as the lack of explicit symmetries like rotations or translations and generally weaker feature correlations compared to the strong spatial or semantic relationships in image or text data [5]. Augmenting tabular data often involves training generative models, which is computationally expensive [51, 12, 25, 50, 26]. On small tasks, this can lead to overfitting [11] and mode collapse [37, 38], where the model fails to capture the diversity of each class. Consequently, existing methods for augmenting tabular data often yield mixed results and can even degrade model performance [27, 40], despite requiring expensive training of generative models.
@@ 3 Method
We propose a general data augmentation framework TabMDA that utilises the in-context learning (ICL) capabilities of any pre-trained tabular in-context model to increase the robustness and performance of any downstream machine learning models (Figure 2). In a nutshell, our approach augments the training dataset by introducing diversity through label-invariant transformations in the manifold space, leveraging multiple contexts for the embedder. We then train the downstream prediction model using this augmented dataset.
@@
In-context Subsetting (ICS). We introduce a manifold data augmentation which performs label-invariant transformations within the manifold (Algorithm 1). The core idea of ICS leverages the context-dependent nature of embeddings. By slightly varying the contexts, ICS generates variations in the embeddings. Specifically, for each data point to be augmented, we generate different contexts by stratified sampling points from the training data without replacement. Each context is denoted as for . TabMDA encodes the data point with each context , producing an augmented embedding . This use of random contexts for encoding enhances data diversity in the manifold space. ICS produces augmented samples for each input.
@@
where and is the dimensionality of the encoder’s latent space. We assume ICS is label-invariant, so all ICS embedding retains the same label as the original sample. The downstream classifier is trained on the augmented dataset , which is times larger than the initial . This augmented dataset is more diverse and includes information from the encoding manifold learned during the pre-training of the in-context encoder. The test data points are encoded using the full training data, which is equivalent to encoding using . As we next demonstrate, TabMDA with ICS improves generalisation and reduces performance variability across classifiers.
@@ 4.1 Experimental Setup
TabMDA settings. For the in-context model, we use the encoder component of TabPFN [17], a pre-trained classifier, without ensembling (referred to as in the original paper). For the proposed in-context subsetting, we define the context size as a proportion of the training set, considering context sizes of 0.5, 0.7, 0.9, and 1. Additionally, we evaluate 5, 20, and 50 subcontexts. We select the optimal TabMDA settings based on the validation balanced accuracy for each dataset configuration.
@@ 4.2 Impact on the Classification Performance
We evaluate the impact of TabMDA on the performance of various tabular classifiers. We consider five classification models with different inductive biases: distance-based methods (K-Nearest Neighbors (KNN)), linear methods (Logistic Regression) and tree-based methods (Decision Tree [6], Random Forest [7] and XGBoost [9]). Following previous work [44], we train the predictors with a fixed set of hyper-parameters, which we provide in Section A.2. We use identical hyper-parameters for the downstream models when trained on real data and with TabMDA.
@@ Acknowledgements
The authors would like to thank Xiangjian Jiang for making the dataset splits. AM acknowledges the support from the Cambridge ESRC Doctoral Training Partnership.
@@ A.2 Implementation, Software and Computing Resources
TabMDA. In all our experiments, TabMDA uses the TabPFN pre-trained in-context model for all experiments, which uses the official weights released by the authors [17] and can be downloaded from https://github.com/automl/TabPFN/raw/main/tabpfn/models_diff/prior_diff_real_checkpoint_n_0_epoch_42.cpkt. We used TabPFN without ensembling (referred to as in the original paper), and employed only its encoder.

### v2

@@ Abstract
Tabular data is prevalent in many critical domains, yet it is often challenging to acquire in large quantities. This scarcity usually results in poor performance of machine learning models on such data. Data augmentation, a common strategy for performance improvement in vision and language tasks, typically underperforms for tabular data due to the lack of explicit symmetries in the input space. To overcome this challenge, we introduce TabMDA, a novel method for manifold data augmentation on tabular data. This method utilises a pre-trained in-context model, such as TabPFN, to map the data into an embedding space. TabMDA performs label-invariant transformations by encoding the data multiple times with varied contexts. This process explores the learned embedding space of the underlying in-context models, thereby enlarging the training dataset. TabMDA is a training-free method, making it applicable to any classifier. We evaluate TabMDA on five standard classifiers and observe significant performance improvements across various tabular datasets. Our results demonstrate that TabMDA provides an effective way to leverage information from pre-trained in-context models to enhance the performance of downstream classifiers. Code is available at https://github.com/AdrianBZG/TabMDA.
@@ 1 Introduction
To address the challenge of small tabular datasets, we introduce TabMDA (Figure 1), a new training-free method for performing manifold data augmentation (MDA) on tabular data. TabMDA is a general method that works with any downstream classifier. It has two components: first, it leverages an existing pre-trained in-context model, such as TabPFN [17], to embed the data in a manifold. Second, we introduce in-context subsetting (ICS), a novel technique that uses an in-context classifier to facilitate label-invariant transformations directly in the manifold space. ICS encodes the input data multiple times using an in-context encoder, with each iteration using different data subsets as context for the transformer model. This process explores the manifold space learned by the pre-trained encoder (see Figure 2 for an illustration). The downstream classifier is then trained on this expanded manifold dataset, thereby indirectly incorporating the pre-training knowledge into any subsequent classification model. Crucially, our method requires no additional training and is applicable to any classifier. We investigate TabMDA using five common tabular classifiers across 28 different dataset configurations of varying sizes. Our results show that TabMDA substantially improves the performance of various standard tabular classifiers, including tree-based ensembles such as XGBoost, while reducing the variability among their performance. As a result, TabMDA enables explainable classifiers, such as KNN, to become very competitive and sometimes even achieve the highest performance.
@@ 2 Related Work
Transformer-based models for tabular data. Transformers for tabular data are typically developed by training from scratch on the target task. Occasionally, an additional pre-training phase on the target task itself is included, unlike LLMs, which use large external corpora. These models introduce various attention mechanisms to handle the tabular data structure. For instance, TUTA [50] introduces a novel attention mechanism that allows the model to attend to the entire table at once, while TAPAS [16] introduces a column-specific attention mechanism that allows the model to attend to the relevant columns in the table. TransTab [51] combines column descriptions and cells as input to a gated Transformer model for feature encoding. SAINT [45] performs attention across both rows and columns to capture the relationships between different cells in the table. More notable examples in this type of models include TabNet [1], TabTransformer [18], and FT-Transformer [14]. Even though transformers can be competitive on large tabular tasks but are often outperformed by well-tuned MLPs [21], especially on small datasets [29]. This suggests that models trained on a single tabular task lack adequate in-context learning (ICL) capabilities for tabular data.
@@
In-context learning (ICL) for tabular data and TabPFN. ICL allows Transformer-based models to adapt to new tasks by conditioning on demonstrations of input-output pairs. Recently, ICL has shown promising results on tabular data. Notably, TabPFN [17], a meta-learned Transformer with ICL capabilities, has demonstrated competitive accuracy compared to tree-based ensembles on small supervised classification tasks. TabPFN employs a Prior-Data Fitted Network (PFN) [33] approach, where it is pre-trained on synthetic datasets generated from Structural Causal Models (SCMs) and Bayesian Neural Networks (BNNs), which resembles the causal structure commonly assumed in tabular data. This pretraining enables TabPFN to approximate Bayesian inference in a single forward pass, reducing computational needs and eliminating hyper-parameter tuning. During inference, TabPFN uses the training dataset (context) and new sample features as input, directly outputting the posterior predictive distribution (PPD) for the new samples. This leverages in-context learning to embed new samples without parameter updates, making TabPFN effective on small datasets. Although TabPFN can embed samples into a rich manifold space using its in-context learning abilities, the nature of this embedding space has not been studied. However, we believe these manifolds could facilitate data augmentation for various classifiers.
@@
Tabular Data Augmentation. Tabular data augmentation is less explored due to several challenges, such as the lack of explicit symmetries like rotations or translations and generally weaker feature correlations compared to the strong spatial or semantic relationships in image or text data [5]. Augmenting tabular data often involves training generative models, which is computationally expensive [53, 12, 25, 52, 26]. On small tasks, this can lead to overfitting [11] and mode collapse [39, 40], where the model fails to capture the diversity of each class. Recently, TabPFGen [27] transformed TabPFN into an energy-based class-conditional generator for data augmentation in the input-space. While it does not require specialised training, TabPFGen achieves minimal performance improvements and retains TabPFN’s limitations, such as being applicable to only up to ten classes. In contrast, our method, TabMDA, is capable of handling datasets with an unlimited number of classes. Overall, existing methods for augmenting tabular data often yield mixed results and can even degrade model performance [28, 42].
@@ 3 Method
We propose a general data augmentation framework TabMDA (Figure 1) that utilises the in-context learning (ICL) capabilities of any pre-trained tabular in-context model to increase the robustness and performance of any downstream machine learning models. In a nutshell, our approach augments the training dataset by introducing diversity through label-invariant transformations in the manifold space, leveraging multiple contexts for the embedder. We then train the downstream prediction model using this augmented dataset.
@@
In-context Subsetting (ICS). We introduce a manifold data augmentation which performs label-invariant transformations within the manifold (Algorithm 1). The core idea of ICS leverages the context-dependent nature of embeddings. By slightly varying the contexts, ICS generates variations in the embeddings (Figure 2). Specifically, for each data point to be augmented, we generate different contexts by stratified sampling points from the training data without replacement. Each context is denoted as for . TabMDA encodes the data point with each context , producing an augmented embedding . This use of random contexts for encoding enhances data diversity in the manifold space. ICS produces augmented samples for each input.
@@
where and is the dimensionality of the encoder’s latent space. We assume ICS is label-invariant, so all ICS embedding retains the same label as the original sample. The downstream classifier is trained exclusively on the augmented dataset , which is the embedding space and is times larger than the initial dataset . This augmented dataset is more diverse and includes information from the encoding manifold learned during the pre-training of the in-context encoder. The test data points are encoded using the full training data, which is equivalent to encoding using . As we next demonstrate, TabMDA with ICS improves generalisation and reduces performance variability across classifiers.
@@ 4.1 Experimental Setup
TabMDA settings. For the in-context model, we use the encoder component of TabPFN [17], a pre-trained classifier. We utilise the embeddings produced by its transformer encoder and we use the official weights released by the TabPFN authors. We use TabPFN without ensembling (referred to as in the original paper), and we do not disable its internal data preprocessing. For the proposed in-context subsetting, we define the context size as a proportion of the training set, considering context sizes of 0.5, 0.7, 0.9, and 1. Additionally, we evaluate 5, 20, and 50 subcontexts. We select the optimal TabMDA settings based on the validation balanced accuracy for each dataset configuration.
@@ 4.2 Impact on the Classification Performance
We assess the impact of TabMDA on the performance of various tabular classifiers. We train these classifiers on real data (the original input space) and on TabMDA-augmented data, which consists of augmented embeddings output from the TabPFN’s transformer encoder. We consider five classification models with different inductive biases: distance-based methods (K-Nearest Neighbors (KNN)), linear methods (Logistic Regression) and tree-based methods (Decision Tree [6], Random Forest [7] and XGBoost [9]). We could not use TabPFN for the final prediction because it handles data with up to 100 dimensions, and the embeddings provided by TabMDA have 512 dimensions. Following previous work [46], we train the predictors with a fixed set of hyper-parameters, which we provide in Section A.2. We use identical hyper-parameters for the downstream models when trained on real data and with TabMDA.
@@ Acknowledgements
The authors would like to thank Xiangjian Jiang for making the dataset splits. AM acknowledges the support from the Cambridge ESRC Doctoral Training Partnership. NS acknowledges the support of the U.S. Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713. Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.